Sandra’s Hair Salon (Q1-Q7) Sandra’s Hair Salon is a hair salon whose service process includes five activities conducted in the sequence described below. (The time required for each activity is shown in parentheses): Activity 1: Welcome a guest and offer homemade herb tea. (10 minutes) Activity 2: Wash and condition hair. (10 minutes) Activity 3: Neck, shoulder, and back stress release massage. (10 minutes) Activity 4: Design the hair style and do the hair. (25 minutes) Activity 5: Check out the guest. (5 minutes) Three servers (S1, S2, and S3) offer the services in a worker-paced line. The assignment of tasks to servers is the following: S1 does Activity 1. S2 does Activities 2 and Activity 3. S3 does Activities 4 and Activity 5. Q1) If the demand of Sandra’s Hair Salon is one customer per hour on Monday mornings, what is the average labor utilization (in %) of the servers?
Answer #1

The data in the question:-

Activity 1: Welcome a guest and offer homemade herb tea. (10 minutes)

Activity 2: Wash and condition hair. (10 minutes)

Activity 3: Neck, shoulder, and back stress release massage. (10 minutes)

Activity 4: Design the hairstyle and do the hair. (25 minutes)

Activity 5: Check out the guest. (5 minutes)

Three servers (S1, S2, and S3) offer the services in a worker-paced line.

The assignment of tasks to servers is the following:

S1 does Activity 1.

S2 does Activities 2 and Activity 3.

S3 does Activities 4 and Activity 5.

Cycle time will be given as below:-

Cycle time =1/(total number of customers /hour) = 1/1 = 1 hour = 60 min/customer.

Total labor used in different tasks:-

S1 does Activity 1:- 10 min.

S2 does Activities 2 and Activity 3:- 10+10=20 min

S3 does Activities 4 and Activity 5:- 25+5=30 min

So total labour content = 10min+20min+30min=60 min

Idle time for S1= total cycle time – S1 processing time =60-10=50 min/customer

Idle time for S2= total cycle time – S2 processing time =60-20=40 min/customer

Idle time for S3= total cycle time – S3 processing time =60-30=30 min/customer

So the total idle time of all the three tasks = (50+40+30)min=120 min/customer

Labour utilization = Total labour content/( total labour content +Total idle time of all the tasks)

Labour utilization =60/(60+120) =60/180 =0.3333 =33.33%

So the answer is:-

Labour utilization =33.33%

KINDLY RATE THE ANSWER.